Guardian Inc. is trying to develop an asset-financing plan. The firm has $330,000 in temporary current assets and $230,000 in permanent current assets. Guardian also has $430,000 in fixed assets. Assume a tax rate of 40 percent.

a. Construct two alternative financing plans for Guardian. One of the plans should be conservative, with 80 percent of assets financed by long-term sources, and the other should be aggressive, with only 56.25 percent of assets financed by long-term sources. The current interest rate is 12 percent on long-term funds and 7 percent on short-term financing. Compute the annual interest payments under each plan.

Annual Interest
Conservative
Aggressive
Answer #1

a. Computation of Annual interest payments

Total Assets = Temporary Assets + Permanent Assets + Fixed Assets

Total Assets = 330000 + 230000 + 430000

Total Assets = $990000

Conservative:

Annual Interest payment = Total Assets * % of Long Term Fianncing * Interest Rate on long term funds + Total Assets * % of Short term financing * Interest rate on short term funds

Annual Interest payment = 990000 * 80% * 12% + 990000 * 20% * 7%

Annual Interest payment = $108900

Aggressive:

Annual Interest payment = Total Assets * % of Long Term Fianncing * Interest Rate on long term funds + Total Assets * % of Short term financing * Interest rate on short term funds

Annual Interest payment = 990000 * 56.25% * 12% + 990000 * 43.75% * 7%

Annual Interest payment = $97143.75